OpenStudy (anonymous):

If the perimeter of a rectangular property is 46 meters, and the area of the property is 76 meters, what is the length of each of the shorter sides?

OpenStudy (anonymous):

2x + 2y =46, and x * Y = 76 2 (76/y) + 2y = 46 The roots are 4 and 19, the smallest is 4

OpenStudy (anonymous):

{2 L + 2 w = 46, L w = 76} {L = 19, w = 4}